AnimeGhoul8863:

Which of the following statements have the same result? Explain each step in solving each one. f(3) when f(x) = 2x + 2 f−1(4) when f(x) = 3x minus 4, all over 5 y + 10 = 2y + 1

dude:

f(3) when f(x) = 2x + 2 f−1(4) when f(x) = 3x minus 4, all over 5 You should probably solve them to have an idea xD

dude:

Do you know what this `f(3) when f(x) = 2x + 2` refers to?

AnimeGhoul8863:

sorry ive been solving it by my self i think i got this but would you check my answers? this is what i have so far

dude:

Okay sure

AnimeGhoul8863:

alright this is my over all answer

AnimeGhoul8863:

I. find f(3) we are going to replace x with 3 so it would look like this f(3)=2(3)+2 once we have this we multiply 2*3 which equals 6 and we plug that in. (3) = 6 + 2 then we add 6+2 which equals 8 which is our answer f(3) = 8 II.find f^-1(4) we are going to replace f(x) with y once we have this we are going y = (3x - 4)/5 to take our equation and swap the variables putting x on the outside x = (3y - 4)/5 and y on the inside next we have to solve for y. which equals 5x = 3y - 4 next we swap sides and place 5x -4 on the left side of the = sign then 5x + 4 = 3y turn the - into a + turn it into a fraction by placing (5x+4) over 3 (5x + 4) over 3 = y f^-1(x) = (5x + 4) over 3 Once you have (5x+4) over 3 we are going to take f^-1(x) f(4) = (5(4) + 4)/3 and replace it with f(4) we then add 4 on both sides and solve f(4) = (20 + 4)/3 this equals 5*4=20 then 20+4=24 which makes the fraction 24/3 f(4) = 24/3 we then divide 24 by 3 which equals 8 f(4) = 8 III. y + 10 = 2y + 13 we are going to start by subtracting 10 from both sides y+10-10=2y+1-10 then we simplify y=2y-9 we then subtract 2y from both sides y-2y=2y-9-2y we then simplify again -y=-9 then we divide by -1 -y -9


=
-1 -1 we simplify one more time and we get our answer 9 this means I and II have the same result of 8.

dude:

You don't have to swap the sides mid step but if it helps then yes Looks good though

AnimeGhoul8863:

Thank you dude im actually proud of my self with how this turned out you guys are alot of help here :D